OpenStudy (anonymous):

Simplify the expression. √6(√33+7) A.√39+7√6 B. √198+7 C, 3√22 D.√198+7√6

sam (.sam.):

√6(√33+7) √6(√33)+√6(7)

sam (.sam.):

thee's one more step to do here, think back, \[\sqrt{a}\sqrt{b}=\sqrt{ab}\]

OpenStudy (anonymous):

B ?

sam (.sam.):

no, try again

OpenStudy (nali):

look it is like this \[\sqrt{6 \times 33}\] + 7\[\sqrt{7}\]

sam (.sam.):

√6(√33)=√(6x33)=√198

OpenStudy (nali):

answer D